# Correlated Proxies: A New Definition and Improved Mitigation for Reward Hacking

### v3

@@ A.2 Failure of Action Distribution Regularization
As discussed in the main text, the OM regularization method we propose differs from the AD-based regularization found in previous work on RLHF. The following theorem shows that almost any form of policy optimization with action distribution regularization cannot guarantee an improvement in true reward over the base policy.
@@
Fix $r\in(0,1)$. Consider a policy optimization objective regularized by any $f$-divergence between the action distributions of the learned policy and the base policy:
$\text{maximize}\qquad L^{\prime}(\pi)=J(\pi,\tilde{R})-J({\pi_{\text{base}}},\tilde{R})-g\left((1-\gamma)\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}D_{f}\Big(\pi(\cdot\mid s_{t})\;\|\;{\pi_{\text{base}}}(\cdot\mid s_{t})\Big)\right]\right),$
@@
Then there is an MDP, reward functions $R,\tilde{R}$, and base policy $\pi_{\text{base}}$} such that $\tilde{R}$ is an $r$-correlated proxy for $R$, but there is a policy $\widetilde{\pi}$ such that
$L^{\prime}(\widetilde{\pi})>0\qquad\text{but}\qquad J(\widetilde{\pi},R)<J({\pi_{\text{base}}},R).$
Theorem A.5 concerns a general form of policy optimization with action distribution regularization; it considers any $f$-divergence between action distributions and any way of scaling the $f$-divergence using the function $g$. For instnce, $g$ could incorporate linear scaling of the divergence as in the KL regularization used in previous work, square-root scaling as we use with $\chi^{2}$ divergence in Theorem 5.1, or any other scaling. The theorem shows that no matter how the regularization is formulated, it cannot guarantee improvement in true reward over the base policy; there is a policy that increases the regularized objective but decreases the true reward.
@@
We consider the base policy defined by
${\pi_{\text{base}}}(a_{1}\mid s_{1})$ $=1\quad$ ${\pi_{\text{base}}}(a_{2}\mid s_{1})$ $=0$
${\pi_{\text{base}}}(a_{1}\mid s_{2})$ $=1\quad$ ${\pi_{\text{base}}}(a_{2}\mid s_{2})$ $=0$
${\pi_{\text{base}}}(a_{1}\mid s_{3})$ $=\gamma\quad$ ${\pi_{\text{base}}}(a_{2}\mid s_{3})$ $=1-\gamma$
${\pi_{\text{base}}}(a_{1}\mid s_{4})$ $=1\quad$ ${\pi_{\text{base}}}(a_{2}\mid s_{4})$ $=0.$
To compute the occupancy measure of the base policy, we first can see that clearly,
$\mu_{{\pi_{\text{base}}}}(s_{1},a_{1})=\mu_{{\pi_{\text{base}}}}(s_{2},a_{1})=\frac{1+r}{4}.$
@@
$\mu_{{\pi_{\text{base}}}}(s_{3})$ $=(1-\gamma)\mu_{0}(s_{3})\left[1+\gamma^{2}+\gamma^{4}+\cdots\right]$
@@
This also implies that $\mu_{{\pi_{\text{base}}}}(s_{4})=\frac{1-r}{4}$ since the occupancy measure sums to one. Thus, we can compute
@@
$\mathbb{E}_{\mu_{{\pi_{\text{base}}}}}[R(s,a)\tilde{R}(s,a)]$ $=r.$
@@
${\pi_{\text{base}}}(a_{1}\mid s_{1})$ $=1\quad$ ${\pi_{\text{base}}}(a_{2}\mid s_{1})$ $=0$
${\pi_{\text{base}}}(a_{1}\mid s_{2})$ $=1\quad$ ${\pi_{\text{base}}}(a_{2}\mid s_{2})$ $=0$
${\pi_{\text{base}}}(a_{1}\mid s_{3})$ $=2\gamma-1\quad$ ${\pi_{\text{base}}}(a_{2}\mid s_{3})$ $=2(1-\gamma)$
${\pi_{\text{base}}}(a_{1}\mid s_{4})$ $=1\quad$ ${\pi_{\text{base}}}(a_{2}\mid s_{4})$ $=0.$
Note that since $\gamma\geq 1/2$ by definition, the policy is well-defined. As for $\pi_{\text{base}}$}, the occupancy measure of $\widetilde{\pi}$ at $s_{1}$ and $s_{2}$ is $\frac{1+r}{4}$. For $s_{3}$, we have
@@
$\leq 0=J({\pi_{\text{base}}},R).$
This verifies the claim that $J(\widetilde{\pi},R)<J({\pi_{\text{base}}},R)$.
@@
$g\left((1-\gamma)\mathbb{E}_{\widetilde{\pi}}\left[\sum_{t=0}^{\infty}\gamma^{t}D_{f}\Big(\widetilde{\pi}(\cdot\mid s_{t})\;\|\;{\pi_{\text{base}}}(\cdot\mid s_{t})\Big)\right]\right)$
$\quad=g\left(\mu(s_{3})D_{f}\Big(\widetilde{\pi}(\cdot\mid s_{3})\;\|\;{\pi_{\text{base}}}(\cdot\mid s_{3})\Big)\right)$
since $\pi_{\text{base}}$} and $\widetilde{\pi}$ only differ in state $s_{3}$. We can rewrite the above as
$=g\left(\frac{(1-r)(1+\gamma)}{4(1+2\gamma)}\left[{\pi_{\text{base}}}(a_{1}\mid s_{3})f\left(\frac{\widetilde{\pi}(a_{1}\mid s_{3})}{{\pi_{\text{base}}}(a_{1}\mid s_{3})}\right)+{\pi_{\text{base}}}(a_{2}\mid s_{3})f\left(\frac{\widetilde{\pi}(a_{2}\mid s_{3})}{{\pi_{\text{base}}}(a_{2}\mid s_{3})}\right)\right]\right)$
@@
$L^{\prime}(\pi)$ $=J(\pi,\tilde{R})-J({\pi_{\text{base}}},\tilde{R})-g\left((1-\gamma)\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}D_{f}\Big(\pi(\cdot\mid s_{t})\;\|\;{\pi_{\text{base}}}(\cdot\mid s_{t})\Big)\right]\right)$
@@ A.4 Learned Reward Functions are $r$-Correlated
Let $R$ be the true reward function and $\tilde{R}$ be a learned reward function. Suppose that $\mathbb{E}_{\mu_{{\pi_{\text{base}}}}}\left[(R(s,a)-\tilde{R}(s,a))^{2}\right]\leq\epsilon\sigma_{R}^{2}$. Then, the learned reward function is an $r$-correlated proxy with $r\geq 1-\epsilon$.
The assumption in Lemma A.8 is that the mean-squared error over the occupancy measure of the base policy is small. This can be achieved, for example, by learning $\tilde{R}$ via least-squares regression over a training dataset of state-action pairs sampled from the base policy. Many results in learning theory show that this results in a small mean-squared error over the distribution the training data was sampled from, i.e., exactly the assumption in Lemma A.8 (Koltchinskii, 2006).
@@
Throughout the proof, all expectations, variances, and covariances are with respect to $\mu_{{\pi_{\text{base}}}}$. We can rewrite the assumption using the bias-variance decomposition as
@@ Appendix B Derivation of Occupancy-Regularized Policy Optimization
$J(\pi_{\theta},\tilde{R})-\lambda\sqrt{\chi^{2}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{base}}}\right)}.$
@@
$\nabla_{\theta}\Big(J(\pi_{\theta},\tilde{R})-\lambda\sqrt{\chi^{2}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{base}}}\right)}\Big)$
$\quad=\nabla_{\theta}J(\pi_{\theta},\tilde{R})-\frac{\lambda\nabla_{\theta}\chi^{2}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{base}}}\right)}{2\sqrt{\chi^{2}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{base}}}\right)}}$
$\quad=\nabla_{\theta}\left(\sum_{s,a}\mu_{\pi_{\theta}}(s,a)\tilde{R}(s,a)\right)-\frac{\lambda}{2\sqrt{\chi^{2}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{base}}}\right)}}\nabla_{\theta}\left(\sum_{s,a}\frac{\mu_{\pi_{\theta}}(s,a)^{2}}{\mu_{{\pi_{\text{base}}}}(s,a)}-1\right)$
$\quad=\sum_{s,a}\left[\nabla_{\theta}\mu_{\pi_{\theta}}(s,a)\tilde{R}(s,a)-\frac{\lambda}{2\sqrt{\chi^{2}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{base}}}\right)}}\nabla_{\theta}\left(\frac{\mu_{\pi_{\theta}}(s,a)^{2}}{\mu_{{\pi_{\text{base}}}}(s,a)}-1\right)\right]$
$\quad=\sum_{s,a}\left[\nabla_{\theta}\mu_{\pi_{\theta}}(s,a)\tilde{R}(s,a)-\frac{\lambda}{2\sqrt{\chi^{2}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{base}}}\right)}}\frac{2\mu_{\pi_{\theta}}(s,a)}{\mu_{{\pi_{\text{base}}}}(s,a)}\nabla_{\theta}\mu_{\pi_{\theta}}(s,a)\right]$
$\quad=\sum_{s,a}\Big(\nabla_{\theta}\mu_{\pi_{\theta}}(s,a)\Big)\left(\tilde{R}(s,a)-\frac{\lambda}{\sqrt{\chi^{2}(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{base}}})}}\frac{\mu_{\pi_{\theta}}(s,a)}{\mu_{\pi_{\text{base}}}(s,a)}\right).$
@@
$R^{\prime}(s,a)=\tilde{R}(s,a)-\frac{\lambda}{\sqrt{\chi^{2}(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{base}}})}}\frac{\mu_{\pi_{\theta}}(s,a)}{\mu_{\pi_{\text{base}}}(s,a)}.$ (19)
However, two terms in (19) cannot be computed directly: the current $\chi^{2}$ divergence between occupancy measures, and the ratio of the occupancy measures $\mu_{\pi_{\theta}}(s,a)/\mu_{\pi_{\text{base}}}(s,a)$. We first show how to approximate the latter using a discriminator network $\hat{d}_{\phi}(s,a)$, trained to optimize
$\phi=\arg\min_{\phi}\;\mathbb{E}_{\mu_{\pi_{\theta}}}\Big[\,\log(1+e^{-\hat{d}_{\phi}(s,a)})\,\Big]\\ \textstyle+\,\mathbb{E}_{\mu_{\pi_{\text{base}}}}\Big[\,\log(1+e^{\hat{d}_{\phi}(s,a)})\,\Big].$ (20)
@@
$\hat{d}_{\phi}(s,a)=\log\frac{\mu_{\pi_{\theta}}(s,a)}{\mu_{\pi_{\text{base}}}(s,a)}.$ (21)
@@
$\chi^{2}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{base}}}\right)=\mathbb{E}_{\mu_{\pi_{\theta}}}\left[\frac{\mu_{\pi_{\theta}}(s,a)}{\mu_{\pi_{\text{base}}}(s,a)}-1\right]=\mathbb{E}_{\mu_{\pi_{\theta}}}\left[e^{\hat{d}_{\phi}(s,a)}-1\right]\eqqcolon\widehat{\chi^{2}}$. (22)
@@
3:  Collect a set of $n$ trajectories $\mathcal{D}_{\pi_{\text{base}}}$ from $\pi_{\text{base}}$}.
@@
$\mathbb{E}_{\mathcal{D}_{\pi}}\Big[\,\log(1+e^{-\hat{d}_{\phi}(s,a)})\,\Big]+\,\mathbb{E}_{\mathcal{D}_{\pi_{\text{base}}}}\Big[\,\log(1+e^{\hat{d}_{\phi}(s,a)})\,\Big]$
@@
$J(\pi_{\theta},\tilde{R})-\lambda\,D_{\text{KL}}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{base}}}\right),$
@@
Figure 6: The top plot shows the state occupancy measures of the base policy $\pi_{\text{base}}$} and a policy $\pi$ optimized with ORPO in the glucose environment. The bottom plot shows the exact log ratio of the occupancy measures $\log\mu_{\pi}(s)/\mu_{\pi_{\text{base}}}(s)$ versus the discriminator output $\hat{d}_{\phi}(s)$, which attempts to approximate it. We find the discriminator output to be a good approximation of the log ratio.

### v4

@@ A.2 Failure of Action Distribution Regularization
As discussed in the main text, the OM regularization method we propose differs from the AD-based regularization found in previous work on RLHF. The following theorem shows that almost any form of policy optimization with action distribution regularization cannot guarantee an improvement in true reward over the reference policy.
@@
Fix $r\in(0,1)$. Consider a policy optimization objective regularized by any $f$-divergence between the action distributions of the learned policy and the reference policy:
$\text{maximize}\qquad L^{\prime}(\pi)=J(\pi,\tilde{R})-J({\pi_{\text{ref}}},\tilde{R})-g\left((1-\gamma)\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}D_{f}\Big(\pi(\cdot\mid s_{t})\;\|\;{\pi_{\text{ref}}}(\cdot\mid s_{t})\Big)\right]\right),$
@@
Then there is an MDP, reward functions $R,\tilde{R}$, and reference policy $\pi_{\text{ref}}$} such that $\tilde{R}$ is an $r$-correlated proxy for $R$, but there is a policy $\widetilde{\pi}$ such that
$L^{\prime}(\widetilde{\pi})>0\qquad\text{but}\qquad J(\widetilde{\pi},R)<J({\pi_{\text{ref}}},R).$
Theorem A.5 concerns a general form of policy optimization with action distribution regularization; it considers any $f$-divergence between action distributions and any way of scaling the $f$-divergence using the function $g$. For instnce, $g$ could incorporate linear scaling of the divergence as in the KL regularization used in previous work, square-root scaling as we use with $\chi^{2}$ divergence in Theorem 5.1, or any other scaling. The theorem shows that no matter how the regularization is formulated, it cannot guarantee improvement in true reward over the reference policy; there is a policy that increases the regularized objective but decreases the true reward.
@@
We consider the reference policy defined by
${\pi_{\text{ref}}}(a_{1}\mid s_{1})$ $=1\quad$ ${\pi_{\text{ref}}}(a_{2}\mid s_{1})$ $=0$
${\pi_{\text{ref}}}(a_{1}\mid s_{2})$ $=1\quad$ ${\pi_{\text{ref}}}(a_{2}\mid s_{2})$ $=0$
${\pi_{\text{ref}}}(a_{1}\mid s_{3})$ $=\gamma\quad$ ${\pi_{\text{ref}}}(a_{2}\mid s_{3})$ $=1-\gamma$
${\pi_{\text{ref}}}(a_{1}\mid s_{4})$ $=1\quad$ ${\pi_{\text{ref}}}(a_{2}\mid s_{4})$ $=0.$
To compute the occupancy measure of the reference policy, we first can see that clearly,
$\mu_{{\pi_{\text{ref}}}}(s_{1},a_{1})=\mu_{{\pi_{\text{ref}}}}(s_{2},a_{1})=\frac{1+r}{4}.$
@@
$\mu_{{\pi_{\text{ref}}}}(s_{3})$ $=(1-\gamma)\mu_{0}(s_{3})\left[1+\gamma^{2}+\gamma^{4}+\cdots\right]$
@@
This also implies that $\mu_{{\pi_{\text{ref}}}}(s_{4})=\frac{1-r}{4}$ since the occupancy measure sums to one. Thus, we can compute
@@
$\mathbb{E}_{\mu_{{\pi_{\text{ref}}}}}[R(s,a)\tilde{R}(s,a)]$ $=r.$
@@
${\pi_{\text{ref}}}(a_{1}\mid s_{1})$ $=1\quad$ ${\pi_{\text{ref}}}(a_{2}\mid s_{1})$ $=0$
${\pi_{\text{ref}}}(a_{1}\mid s_{2})$ $=1\quad$ ${\pi_{\text{ref}}}(a_{2}\mid s_{2})$ $=0$
${\pi_{\text{ref}}}(a_{1}\mid s_{3})$ $=2\gamma-1\quad$ ${\pi_{\text{ref}}}(a_{2}\mid s_{3})$ $=2(1-\gamma)$
${\pi_{\text{ref}}}(a_{1}\mid s_{4})$ $=1\quad$ ${\pi_{\text{ref}}}(a_{2}\mid s_{4})$ $=0.$
Note that since $\gamma\geq 1/2$ by definition, the policy is well-defined. As for $\pi_{\text{ref}}$}, the occupancy measure of $\widetilde{\pi}$ at $s_{1}$ and $s_{2}$ is $\frac{1+r}{4}$. For $s_{3}$, we have
@@
$\leq 0=J({\pi_{\text{ref}}},R).$
This verifies the claim that $J(\widetilde{\pi},R)<J({\pi_{\text{ref}}},R)$.
@@
$g\left((1-\gamma)\mathbb{E}_{\widetilde{\pi}}\left[\sum_{t=0}^{\infty}\gamma^{t}D_{f}\Big(\widetilde{\pi}(\cdot\mid s_{t})\;\|\;{\pi_{\text{ref}}}(\cdot\mid s_{t})\Big)\right]\right)$
$\quad=g\left(\mu(s_{3})D_{f}\Big(\widetilde{\pi}(\cdot\mid s_{3})\;\|\;{\pi_{\text{ref}}}(\cdot\mid s_{3})\Big)\right)$
since $\pi_{\text{ref}}$} and $\widetilde{\pi}$ only differ in state $s_{3}$. We can rewrite the above as
$=g\left(\frac{(1-r)(1+\gamma)}{4(1+2\gamma)}\left[{\pi_{\text{ref}}}(a_{1}\mid s_{3})f\left(\frac{\widetilde{\pi}(a_{1}\mid s_{3})}{{\pi_{\text{ref}}}(a_{1}\mid s_{3})}\right)+{\pi_{\text{ref}}}(a_{2}\mid s_{3})f\left(\frac{\widetilde{\pi}(a_{2}\mid s_{3})}{{\pi_{\text{ref}}}(a_{2}\mid s_{3})}\right)\right]\right)$
@@
$L^{\prime}(\pi)$ $=J(\pi,\tilde{R})-J({\pi_{\text{ref}}},\tilde{R})-g\left((1-\gamma)\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}D_{f}\Big(\pi(\cdot\mid s_{t})\;\|\;{\pi_{\text{ref}}}(\cdot\mid s_{t})\Big)\right]\right)$
@@ A.4 Learned Reward Functions are $r$-Correlated
Let $R$ be the true reward function and $\tilde{R}$ be a learned reward function. Suppose that $\mathbb{E}_{\mu_{{\pi_{\text{ref}}}}}\left[(R(s,a)-\tilde{R}(s,a))^{2}\right]\leq\epsilon\sigma_{R}^{2}$. Then, the learned reward function is an $r$-correlated proxy with $r\geq 1-\epsilon$.
The assumption in Lemma A.8 is that the mean-squared error over the occupancy measure of the reference policy is small. This can be achieved, for example, by learning $\tilde{R}$ via least-squares regression over a training dataset of state-action pairs sampled from the reference policy. Many results in learning theory show that this results in a small mean-squared error over the distribution the training data was sampled from, i.e., exactly the assumption in Lemma A.8 (Koltchinskii, 2006).
@@
Throughout the proof, all expectations, variances, and covariances are with respect to $\mu_{{\pi_{\text{ref}}}}$. We can rewrite the assumption using the bias-variance decomposition as
@@ Appendix B Derivation of Occupancy-Regularized Policy Optimization
$J(\pi_{\theta},\tilde{R})-\lambda\sqrt{\chi^{2}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{ref}}}\right)}.$
@@
$\nabla_{\theta}\Big(J(\pi_{\theta},\tilde{R})-\lambda\sqrt{\chi^{2}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{ref}}}\right)}\Big)$
$\quad=\nabla_{\theta}J(\pi_{\theta},\tilde{R})-\frac{\lambda\nabla_{\theta}\chi^{2}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{ref}}}\right)}{2\sqrt{\chi^{2}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{ref}}}\right)}}$
$\quad=\nabla_{\theta}\left(\sum_{s,a}\mu_{\pi_{\theta}}(s,a)\tilde{R}(s,a)\right)-\frac{\lambda}{2\sqrt{\chi^{2}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{ref}}}\right)}}\nabla_{\theta}\left(\sum_{s,a}\frac{\mu_{\pi_{\theta}}(s,a)^{2}}{\mu_{{\pi_{\text{ref}}}}(s,a)}-1\right)$
$\quad=\sum_{s,a}\left[\nabla_{\theta}\mu_{\pi_{\theta}}(s,a)\tilde{R}(s,a)-\frac{\lambda}{2\sqrt{\chi^{2}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{ref}}}\right)}}\nabla_{\theta}\left(\frac{\mu_{\pi_{\theta}}(s,a)^{2}}{\mu_{{\pi_{\text{ref}}}}(s,a)}-1\right)\right]$
$\quad=\sum_{s,a}\left[\nabla_{\theta}\mu_{\pi_{\theta}}(s,a)\tilde{R}(s,a)-\frac{\lambda}{2\sqrt{\chi^{2}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{ref}}}\right)}}\frac{2\mu_{\pi_{\theta}}(s,a)}{\mu_{{\pi_{\text{ref}}}}(s,a)}\nabla_{\theta}\mu_{\pi_{\theta}}(s,a)\right]$
$\quad=\sum_{s,a}\Big(\nabla_{\theta}\mu_{\pi_{\theta}}(s,a)\Big)\left(\tilde{R}(s,a)-\frac{\lambda}{\sqrt{\chi^{2}(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{ref}}})}}\frac{\mu_{\pi_{\theta}}(s,a)}{\mu_{\pi_{\text{ref}}}(s,a)}\right).$
@@
$R^{\prime}(s,a)=\tilde{R}(s,a)-\frac{\lambda}{\sqrt{\chi^{2}(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{ref}}})}}\frac{\mu_{\pi_{\theta}}(s,a)}{\mu_{\pi_{\text{ref}}}(s,a)}.$ (19)
However, two terms in (19) cannot be computed directly: the current $\chi^{2}$ divergence between occupancy measures, and the ratio of the occupancy measures $\mu_{\pi_{\theta}}(s,a)/\mu_{\pi_{\text{ref}}}(s,a)$. We first show how to approximate the latter using a discriminator network $\hat{d}_{\phi}(s,a)$, trained to optimize
$\phi=\arg\min_{\phi}\;\mathbb{E}_{\mu_{\pi_{\theta}}}\Big[\,\log(1+e^{-\hat{d}_{\phi}(s,a)})\,\Big]\\ \textstyle+\,\mathbb{E}_{\mu_{\pi_{\text{ref}}}}\Big[\,\log(1+e^{\hat{d}_{\phi}(s,a)})\,\Big].$ (20)
@@
$\hat{d}_{\phi}(s,a)=\log\frac{\mu_{\pi_{\theta}}(s,a)}{\mu_{\pi_{\text{ref}}}(s,a)}.$ (21)
@@
$\chi^{2}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{ref}}}\right)=\mathbb{E}_{\mu_{\pi_{\theta}}}\left[\frac{\mu_{\pi_{\theta}}(s,a)}{\mu_{\pi_{\text{ref}}}(s,a)}-1\right]=\mathbb{E}_{\mu_{\pi_{\theta}}}\left[e^{\hat{d}_{\phi}(s,a)}-1\right]\eqqcolon\widehat{\chi^{2}}$. (22)
@@
3:  Collect a set of $n$ trajectories $\mathcal{D}_{\pi_{\text{ref}}}$ from $\pi_{\text{ref}}$}.
@@
$\mathbb{E}_{\mathcal{D}_{\pi}}\Big[\,\log(1+e^{-\hat{d}_{\phi}(s,a)})\,\Big]+\,\mathbb{E}_{\mathcal{D}_{\pi_{\text{ref}}}}\Big[\,\log(1+e^{\hat{d}_{\phi}(s,a)})\,\Big]$
@@
$J(\pi_{\theta},\tilde{R})-\lambda\,D_{\text{KL}}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{ref}}}\right),$
@@
Figure 6: The top plot shows the state occupancy measures of the reference policy $\pi_{\text{ref}}$} and a policy $\pi$ optimized with ORPO in the glucose environment. The bottom plot shows the exact log ratio of the occupancy measures $\log\mu_{\pi}(s)/\mu_{\pi_{\text{ref}}}(s)$ versus the discriminator output $\hat{d}_{\phi}(s)$, which attempts to approximate it. We find the discriminator output to be a good approximation of the log ratio.